1,3,5,7,9
What is the sum of all 5 digit no formed by above given no. repetition is not allowed

Answer by ikleyn(52932)

.
  (1+3+5+7+9) +

+ (1+3+5+7+9)*10 +

+ (1+3+5+7+9)*100 +

+ (1+3+5+7+9)*1000 +

+ (1+3+5+7+9)*10000    = 25*11111 = 277775.

Solved.